2 parts thx :)

1. A point source emits sound waves isotropically. A sound meter measures a sound level of 51 dB at location C and a sound level of 48.47 at location D, a distance of 16 m from location C. The two locations C and D and the point source are all located along the same line. How far from location C is the point source? Give your answer in m, though enter only the numerical part in the box.

2. A point source emits sound waves isotropically. A sound meter measures a sound level of 51.08 dB at location C and a sound level of 45.52 at location D, a distance of 15 m from location C. The two locations C and D and the point source are all located along the same line. What is the total output power emitted by the point source? Give your answer in mW, though enter only the numerical part in the box.
Answer #1

Intensity(I) is inversely proportional to the square of the distance from source(r2)

  Ia(1/2)

a)di 51dB to=48:47 Gource k a h 16m let sound source intensity I, a RT = L. · (Intensity is in veysely propertienal to - qquarb)I -51.05 Ch 10 = 45.52 dB. source 4 sm ism source I ago 72 51.03 t1572 Ic = 7+15) 45.52 spre Io 1+ 15 = 1.0593 15 = 0.0593 x